OpenStudy (anonymous):

Please help me with this: √c (3+2√5)

OpenStudy (anonymous):

@jim_thompson5910 @e.cociuba

OpenStudy (anonymous):

@divu.mkr @shrutipande9 @RadEn Can you help my friend?

OpenStudy (anonymous):

@surjithayer can you help my friend

OpenStudy (amistre64):

the question is too vague

OpenStudy (anonymous):

@amistre64 what do u mean?

OpenStudy (anonymous):

what is wanted in the question is not clear.

OpenStudy (anonymous):

@surjithayer to simplify it

OpenStudy (anonymous):

\[\sqrt{c \left( 3+2\sqrt{5} \right)}=\sqrt{3c+2c \sqrt{5}}\]

OpenStudy (anonymous):

or\[\sqrt{c}\left( 3+2\sqrt{5} \right)=3\sqrt{c}+2\sqrt{5c}\] if radical sign on c only.